An engine using 1 mol of an ideal gas initially at 22.3 L and 454 K performs a cycle consisting of four steps: 1) an isothermal expansion at 454 K from 22.3 L to 40 L ; 2) cooling at constant volume to 270 K ; 3) an isothermal compression to its original volume of 22.3 L; and 4) heating at constant volume to its original temperature of 454 K . Find its efficiency. Assume that the heat capacity is 21 J/K and the universal gas constant is 0.08206 L · atm/mol/K = 8.314 J/mol/K. Answer in units of %.
Answer #1

Ans: Given n = 1 mol, V1 = 22.3 L, T1 = 454 K => P1 = nRT1/V1 = 1.671 atm

Step-1: T = 0 => U = 0, T2 = 454 K, V2 = 40L => P2 = nRT2/V2 = 0.931 atm

Work done by gas

From 1st law of thermodynamics,

Step-2: V = 0 =>V3 = 40L, W = 0, T3 = 270 K => T = 270 - 454 = -184 K

W2 = 0 J

Q2 = U = nCT = 1 x 21 x -184 = -3864 J

Step-3:  T = 0 => U = 0, T4 = 270 K, V4 = 22.3L => P4 = nRT4/V4 = 0.994 atm

Work done by gas

From 1st law of thermodynamics,

Step-4: V = 0 => W = 0 => T = 454-270 = 184 K

W2 = 0 J

Q4 = U = nCT = 1 x 21 x 184 = 3864 J

Total Heat given to the system is Q = Q1 +Q4 = 6069.446 J

Net work done W = W1+ W2+ W3+ W4 = 893.837 J

Efficiency = W/Q = 0.1473 = 14.73%